Carson Electronics' management has long viewed BGT Electronics as an industry leader and uses this firm as a model firm for analyzing its own performance. The balance sheets and income statements for the two firms are found below

a. Calculate the following ratios for both Carson and BGT:

Current ratio

Times interest earned

Inventory turnover

Total asset turnover

Operating profit margin

Operating return on assets

Debt ratio

Average collection period

Fixed asset turnover

Return on equity

b. Analyze the differences you observe between the two firms. Comment on what you view as weaknesses in the performance of Carson as compared to BGT that Carson's management might focus on to improve its operations.

Balance Sheet ($000) Carson Electronics, Inc. BGT Electronics, Inc.
Cash $2,030 $1,450
Accounts receivable 4470 6020
Inventories 1490 2530
Current assets $7,990 $10,000
Net fixed assets 16050 25020
Total assets $24,040 $35,020
Accounts payable $2,450 $5,010
Accrued expenses 1050 1500
Short-term notes payable 3480 1460
Current liabilities $6,980 $7,970
Long-term debt 8050 3960
Owners' equity 9010 23090
Total liabilities and owners' equity $24,040 $35,020
Income Statement ($000) Carson Electronics, Inc. BGT Electronics, Inc.
Net sales (all credit) $47,990 $70,000
Cost of goods sold (36,020) (42,030)
Gross profit $11,970 $27,970
Operating expenses (8,020) (12,010)
Net operating income $3,950 $15,960
Interest expense (1,190) (540)
Earnings before taxes $2,760 $15,420
Income taxes (40%) (1,104) (6,168)
Net income $1,656 $9,252
Answer #1

a) i) Current ratio = Current assets / Current liabilities

Here,

Current assets = Cash + Accounts recievable + Inventories

Current liabilities = Accounts payable + Accrued expenses + Short Term notes payable

Current ratio (Carson) = (2030 + 4470 + 1490) / (2450 + 1050 + 3480) = 7990 / 6980 = 1.15

Current ratio (BGT) = (1450 + 6020 + 2530) / (5010 + 1500 + 1460) = 10000 / 7970 = 1.26

ii) Times interest earned = Earnings before interest & tax (EBIT) / Interest

Here, EBIT means net operating income

Times interest earned (Carson) =3950/1190 = 3.32

Times interest earned (BGT) = 15960/540 = 29.56

iii) Inventory turnover = Cost of goods sold (COGS) / Inventories

Inventory turnover (Carson) = 36020/1490 = 24.18

Inventory turnover (BGT) = 42030 / 2530 = 16.61

iv) Total assets turnover = Sales / Total assets

Total assets turnover (Carson)=47990/24040 = 2.0

Total assets turnover (BGT) = 70000/35020 = 2.0

v) Operating profit margin = Operating profit (EBIT) / Sales * 100

Operating profit margin (Carson) = 3950 / 47990 * 100 = 8.23%

Operating profit margin (BGT) = 15960 / 70000 * 100 = 22.80%

vi) Operating return on assets = Operating return (EBIT) / Total assets * 100

Operating return on assets (Carson) = 3950 / 24040 * 100 = 16.43%

Operating return on assets (BGT) = 15960 / 35020 * 100 = 45.57%

vii) Debt ratio = Total liabilities / Total assets

Here,

Total liabilities = Current liabilities + Long term debt

Debt ratio (Carson) = (6980 + 8050) / 24040 = 0.63

Debt ratio (BGT) = (7970 + 3960) / 35020 = 0.34

viii) Average collection period = Accounts recievables / Sales * 365 days

Average collection period (Carson) = 4470 / 47990 * 365 = 34 days

Average collection period (BGT) = 6020 / 70000 * 365 days = 31.39 days

ix) Fixed assets turnover = Sales / Fixed assets

Fixed assets turnover (Carson) = 47990 / 16050 = 2.99

Fixed assets turnover (BGT) = 70000/25020 = 2.80

x) Return on equity = Profit after tax / Net worth

Here,

Net worth = Equity + Reserves

Return on equity (Carson) = 1656/(9010+0) = 0.18

Return on equity (BGT) = 9252 / (23090+0) = 0.40

b) Reason for weakness in performance of Carson:

i) Times interest earned ratio is much lower than BGT which affects the profit margin

ii) Operating profit margin & return on assets margin is much lower than BGT which shows Carson's inefficiency.

iii) Debt ratio is much higher than BGT which results in higher interest payment and lower of profits

iv) Return on equity is also lower than BGT which is also the reason behind for Carson's weaker performance.